# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 1.3 Comment on the proofs
In summary, this strategy relies on three major components. First is to understand cycles of length $\leq 6$ in $\Delta_{S}$ with $S$ spherical. The second is to prove certain properties of cycles up to a given length (mostly up to length $6$) survive under the propagation process mentioned above. Third, in order to extend this strategy so that it can treat all Artin groups, we need a theorem reducing the non-positive curvature of $\Delta_{S,S^{\prime}}$ to understanding cycles in vertex links up to a given length in $\Delta_{S,S^{\prime}}$ when $S^{\prime}$ belongs to one of the six exceptional cases. Even without the third component, being able to treat the first and second components in full generality will already lead to fairly general results in the $K(\pi,1)$-conjecture.
What is new in this article.
As explained before, a main challenge in implementing the strategy in [Hua23] is to understand the filling of cycles of length $\leq 6$ in the Artin complex $\Delta_{S}$ with the associated Artin group $A_{S}$ being spherical. Almost half of [Hua23] is to understand how $4$-cycles in $\Delta_{S}$ can be filled in the complex. This relies on understanding equation (1.14) in the case $n=4$. The main observation was that solving such an equation can be reduced to understanding when two elements in the group commute. However, this completely breaks down in the case $n>4$, which makes understanding 5-cycles and 6-cycles in $A_{S}$ a different story. This is a major obstacle to generalizing [Hua23].
A big portion of this article (Section 4 to Section 7), as well as the companion article [Hua24] is devoted to developing new methods to understand 6-cycles in some Artin complexes of spherical types, providing strong evidence for overcoming the obstacle in the previous paragraph. The case discussed in the companion article [Hua24] uses methods of different flavors, hence is separated from here. We are able to treat certain types of 6-cycles in $\Delta_{S}$ with $S$ being of type $F_{4},A_{n},D_{n},H_{3}$, where the $D_{n}$ case is in [Hua24].
We introduce a new ingredient in Section 8 on how some of the properties we prove for spherical type Artin complexes propagate to more general Artin complexes in the sense explained before. In Section 9 to Section 11, we use the results from section 4 to Section 8, as well as the method in [Hua23], to prove all the main theorems.  Here we only discuss some of the material in Section 4 to Section 8.
When $A_{S}$ is spherical, the Artin complex $\Delta_{S}$ is also called the spherical Deligne complex. It is also named “generalized building” by Deligne [Del72], due to its similarity with spherical buildings. These complexes are unions of “apartments” that are round spheres, however, they are not exactly spherical buildings. Nevertheless, this analogy is  fruitful and influential, leading Charney and Davis to conjecture that spherical Deligne complexes are CAT$(1)$ with a naturally defined metric, as a way to prove the $K(\pi,1)$-conjecture.
On the other hand, the main property we want to establish, is that given a 6-cycle in the 1-skeleton of $\Delta_{S}$, the minimal filling of this 6-cycle in the 2-skeleton of $\Delta_{S}$ is restricted to a few very specific combinatorial configurations. This is a combinatorial problem that is a reminiscent of Bowditch’s characterization of CAT$(1)$ in terms of filling curves of length $<2\pi$ [Bow95].
These spherical Deligne complexes are more commonly thought of as “spherical objects” and were studied using intuition from spherical geometry. However, in this article, we treat them as “Euclidean objects”. This viewpoint is motivated by certain topological properties of the associated hyperplane arrangement complements. We show these spherical complexes contain large pieces satisfying fine notions of non-positive curvature, and a large amount of group theoretical information is encoded in these pieces. While the meaning of pieces and the notions of non-positive curvature vary in cases, this is the organizing principle for us to understand filling cycles in the spherical Deligne complexes. The choice of the form of non-positive curvature is quite subtle - it was pointed out to the author by M. Bestvina that there are other contexts (see e.g. curve complexes and arc complexes [Web20]) where a given complex satisfies nice form of negative curvature/non-positive curvature, but fails to meet the types of conditions we want to verify.
Now we discuss cases $A_{n},H_{3}$ in more detail on how we find these non-positive curvature subcomplexes, where we actually have control over filling cycles of any length in these subcomplexes. Suppose $A_{S}$ is an Artin group of type $A_{n}$, with the associated pure Artin group denoted by $PA_{S}$. Then $\Delta_{S}/PA_{S}$ is a copy of the Coxeter complex $\mathsf{C}_{S}$. This gives a folding map $\pi:\Delta_{S}\to\mathsf{C}_{S}$. Let $H$ be a wall in $\mathsf{C}_{S}$ (i.e., fix points of a reflection), and let $K_{H}$ be the maximal subcomplex of $\mathsf{C}_{S}$ contained in an open halfspace bounded by $H$. Let $\widetilde{K}_{H}=\pi^{-1}(K_{H})$. While $K_{H}$ has a natural piecewise spherical metric inherited from the round sphere $\mathsf{C}_{S}$, we would like to metrize $K_{H}$ differently.
Consider the Boolean lattice which is the poset made of all subsets of an $n$-element set, ordered under containment. The Hasse diagram of this poset can be identified as the 1-skeleton of an $n$-dimensional unit cube $[0,1]^{n}$.  A simple computation yields that $K_{H}$ is isomorphic (as a simplicial complex) to the geometric realization of this poset, which corresponds to a simplicial subdivision of $[0,1]^{n}$ into orthoschemes. We endow the unit cube with $\ell^{\infty}$-metric, which gives a metric on $K_{H}$, and pulls back to a piecewise $\ell^{\infty}$ metric on $\widetilde{K}_{H}$. The following is the main point, and follows from Proposition 5.4.

### v3

@@ 2.5 Subdivision of some relative Artin complexes
When $\dim(\Delta_{S,S^{\prime}})>2$, the link conditions in Corollary 2.27 amounts to understand certain $n$-cycles in $\Delta_{S,S^{\prime}}$ with $n\leq 6$. More precisely, each bowtie in $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$ gives rise to a 4-cycle in $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$. Checking the bowtie free amounts to check if such 4-cycle is embedded and induced, then it has a center in the sense explained in Section 1.2. See Lemma 2.30 for a precise statement. In particular, each such $4$-cycle can be filled by a disk in the 2-skeleton made of four 2-simplices. To check a flagness condition, say we have vertices $x_{1},x_{2},x_{3}$ in $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$ which are pairwisely upper bounded. Let $y_{i}$ be an upper bound for $x_{i}$ and $x_{i+1}$ for $i\in\mathbb{Z}/3\mathbb{Z}$. Then $x_{1}y_{1}x_{2}y_{2}x_{3}y_{3}$ forms a 6-cycle in $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$. Checking upward flagness amounts to find a vertex $z\in\operatorname{lk}(x,\Delta_{S,S^{\prime}})$ adjacent to each of $\{x_{1},x_{2},x_{3}\}$ which is a common upper bound of them. Note that the new vertex $z$ breaks down the 6-cycle into three 4-cycles, namely $zx_{1}y_{1}x_{2}$, $zx_{2}y_{2}x_{3}$ and $zx_{3}y_{3}x_{1}$ such that each of them can be filled in as in the previous paragraph if we know the bowtie free condition. Thus checking the flagness condition is reduced to showing certain 6-cycles can be filled in the 2-skeleton by disks with specific combinatorial types.
Definition 2.29.
Let $\Lambda$ be a Coxeter diagram which is a tree, with its set of nodes $S$. Let $Z$ be a simplicial complex of type $S$. Let $X$ be the 1-skeleton of $Z$ with its vertex types as explained above. We say $Z$ satisfies the labeled 4-wheel condition if for any induced 4-cycle in $X$ with consecutive vertices being $\{x_{i}\}_{i=1}^{4}$ and their types being $\{\hat{s}_{i}\}_{i=1}^{4}$, there exists a vertex $x\in X$ adjacent to each of $x_{i}$ such that the type $\hat{s}$ of $x$ satisfies that the node $s$ is in the smallest subtree of $\Lambda^{\prime}$ containing all of $\{s_{i}\}_{i=1}^{4}$.
The following is a consequence of [Hua24a, Lemma 6.14 and Proposition 6.17].
Lemma 2.30.
Note that Lemma 2.27 requires $\Lambda^{\prime}$ to be an admissible subgraph of $\Lambda$. If $\Lambda^{\prime}$ is not an admissible subgraph of $\Lambda$, then the link criterion in Lemma 2.27 does not apply directly. For example, in Case 2 of Lemma 2.27, if $\Lambda^{\prime}$ is not admissible, then $\Delta_{S,S^{\prime}}$ with the relation induced from the linear order from $S^{\prime}$ is not a poset. So Assumption 2 of Theorem 2.18 is not satisfied. Similar problems happen with the other cases. This leads to us to formulate the following somewhat less precise conjecture.
Conjecture 2.31.
Under the same assumption of Lemma 2.27, even if $\Lambda^{\prime}$ is not admissible in $\Lambda$, there still exists a criterion only involving cycles in the 1-skeleton in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then $\Delta_{S,S^{\prime}}$ is contractible.
For the remaining types of $\Lambda^{\prime}$, we ask the following.
Question 2.32.
Suppose $\Lambda$ is a connected Coxeter diagram, and $\Lambda^{\prime}$ is an induced subgraph of $\Lambda$ in the sense of Definition 2.16. Suppose the type of $\Lambda^{\prime}$ is one of $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, [3,5,3],[5,3,3,3]}[3,5,3],[5,3,3,3]\}[ 3 , 5 , 3 ] , [ 5 , 3 , 3 , 3 ] }. Is it true that there is a criterion only involving cycles in the 1-skeleton in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then $\Delta_{S,S^{\prime}}$ is contractible?
We firmly believe the answer to this question is yes - this is ongoing work of P. Przytycki and the author, and another ongoing work of K. Goldman and P. Przytycki. This question is of independent interests outside $K(\pi,1)$-conjecture, as it would lead to new form of non-positively curvature which can be possibly applied to other classes of groups or spaces.
2.6 Link conditions at the base case
Let $S^{\prime}\subset S$ as before with $S^{\prime}$ almost spherical. We are finally in the position to discuss checking link conditions on $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$, which is the core of our approach to $K(\pi,1)$-conjecture. Our plan is to reduce the study of cycles in general relative Artin complexes to the base case of studying cycles in Artin complexes associated with spherical Artin groups (i.e. cycles in spherical Deligne complexes), so eventually proving $K(\pi,1)$-conjecture for general Artin groups can be reduced to specific properties of spherical Artin groups.
For a vertex $x$ of type $\hat{s}$ in $\Delta_{S,S^{\prime}}$, by Lemma 2.2, $\operatorname{lk}(x,\Delta_{S,S^{\prime}})\cong\Delta_{T,T^{\prime}}$ with $T=S\setminus\{s\}$ and $T^{\prime}=S^{\prime}\setminus\{s\}$. If $A_{T}$ is spherical, then we are already in the base case. If $A_{T}$ is not spherical, then we can find $T^{\prime\prime}\subset T$ such that $A_{T^{\prime\prime}}$ is almost spherical. If we understand certain collections of cycles in the vertex links of $\Delta_{T,T^{\prime\prime}}$, then we can conclude that $\Delta_{T,T^{\prime\prime}}$ is non-positively curved in an appropriate sense, using some of the criterion in Section 2.4 and Section 2.5. While $T^{\prime\prime}$ might not equal $T^{\prime}$, $A_{T}$ acts on both $A_{T,T^{\prime\prime}}$ and $A_{T,T^{\prime}}$, this allows us to encode an $n$-cycle in $\Delta_{T,T^{\prime}}$ as a collection of convex subspaces of $\Delta_{T,T^{\prime\prime}}$, and use the non-positive curvature of $\Delta_{T,T^{\prime\prime}}$ to analyze the configuration of these subspaces. This gives a way to use our knowledge of cycles in the vertex links of $\Delta_{T,T^{\prime\prime}}$ associated to a smaller Artin group $A_{T}$ to understand cycles in the vertex links of $\Delta_{S,S^{\prime}}$ associated with a larger Artin group $A_{S}$. We can keep doing this until $A_{T}$ is spherical. We refer to [Hua24a, Section 9.1] for an example demonstrating this strategy.
Now we look at the base case $A_{T}=A_{S\setminus\{s\}}$ being spherical in greater detail. By translating the meaning of the link conditions in Corollary 2.27 in the base case, we are led to the following statements about spherical Artin groups.
Conjecture 2.33.
Suppose $A_{S}$ is an irreducible spherical Artin group with Coxeter diagram $\Lambda$. Let $\Lambda^{\prime}$ be a linear subdiagram of $\Lambda$ with consecutive vertices $\{s_{i}\}_{i=1}^{n}$ such that the edge between $s_{n-1}$ and $s_{n}$ has label $\geq 4$. Then the vertex set of $\Delta_{\Lambda,\Lambda^{\prime}}$ equipped with the relation induced from $s_{1}<s_{2}<\cdots<s_{n}$ (as in Definition 2.15) is a bowtie free and upward flag poset.
Conjecture 2.34.
Suppose $A_{S}$ is an irreducible spherical Artin group. Let $S^{\prime}\subset S$ such that $A_{S^{\prime}}$ has Coxeter diagram isomorphic to the type $D_{n}$ Coxeter diagram for $n\geq 3$ (the isomorphism does not need to preserve edge labels). Let $\{b_{i}\}_{i=1}^{n+1}$ be vertices in $S^{\prime}$ as in Figure 3 left. Then the $(b_{1},b_{2})$-subdivision of $\Delta_{S,S^{\prime}}$ (in the sense of Definition 2.21) is a bowtie free and downward flag poset.
In Conjecture 2.34, we allow the case when the Dynkind diagram $A_{S^{\prime}}$ is isomorphic to the type $D_{3}$ diagram (though edge labels might not be preserved). While $D_{3}$ diagram is the same as $A_{3}$ diagram, but we are considering a subdivision of $\Delta_{S,S^{\prime}}$ by viewing the Coxeter diagram as $D_{3}$ rather than $A_{3}$, as explained in Definition 2.21.
Now we summarize previous results on these conjectures. The bowtie free part of both conjectures is already known and is a consequence of Theorem 2.35 and Theorem 2.36.
Theorem 2.35.
([Hua24a, Theorem 8.1]) Suppose $A_{\Lambda}$ is an irreducible spherical Artin group. Then for any linear subgraph $\Lambda^{\prime}\subset\Lambda$, $\Delta_{\Lambda,\Lambda^{\prime}}$ is bowtie free.
This theorem is a consequence of Theorem 2.36 below and Lemma 2.30.
Theorem 2.36.
([Hua24a, Proposition 2.8]) Suppose $A_{S}$ is an irreducible spherical Artin group. Then $\Delta_{S}$ satisfies the labeled 4-wheel condition.
For the upward flag part of Conjecture 2.33, the case when $A_{S}=A_{S^{\prime}}$ is type $B_{n}$ follows from the work of Haettel [Hae21, Proposition 6.6].
Theorem 2.37.
Let $A_{S}$ be the Artin group of type $B_{n}$. Let $S=\{s_{1},s_{2},\ldots,s_{n}\}$ such that $s_{i}$ and $s_{i+1}$ are adjacent in the Coxeter diagram and $m_{s_{n-1},s_{n}}=4$. We put a total order on $S$ by $s_{1}<s_{2}<\ldots<s_{n}$. Let the vertex set $V$ of $\Delta_{S}$ be endowed with the relation $<$ induced from this partial order on $S$. Then $(V,\leq)$ is an upward flag poset.
For Conjecture 2.34, the downward flag part is true when $A_{S}=A_{S^{\prime}}$ is of type $D_{4}$, see [Hua24b, Corollary 7.7]. The downward flag part is also known when $A_{S}$ is of type $D_{n}$, and $S^{\prime}=\{b_{1},b_{2},b_{3}\}$ in Figure 3, see Theorem 2.38 below and Lemma 8.2.
Theorem 2.38.
([Hua24b, Theorem 1.2]) Let $\Lambda$ be the Coxeter diagram of type $D_{n}$. Let $\Lambda^{\prime}\subset\Lambda$ be the subgraph spanned by $\{\delta_{1},\delta_{2},\delta_{3}\}$ in Figure 4. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is weakly flag.
Figure 4: Coxeter diagram of type $D_{n}$.
3 One side flagness in type $F_{4}$
The main goal of this section is Proposition 3.2. The proof is this proposition is inspired by [Hae21, Proposition 6.6].
Lemma 3.1.
Suppose $A_{\Lambda}$ is a spherical Artin group with Coxeter diagram $\Lambda$ and suppose $\alpha:A_{\Lambda}\to A_{\Lambda}$ be an automorphism induced by an automorphism of $\Lambda$. Let $C$ be a left coset of a standard parabolic subgroup of $A_{\Lambda}$. If $\alpha(C)=C$, then $\alpha$ fixes an element in $C$.
Proof.
[…]